Jane invests $9,885 into two accounts. A savings account earns 6% interest , while a money market account earns 9.5% interest. After one year she made $696.35 interest from both accounts . How much did she invest into each account?
Answer by stanbon(75887)

Jane invests $9,885 into two accounts. A savings account earns 6% interest , while a money market account earns 9.5% interest. After one year she made $696.35 interest from both accounts . How much did she invest into each account?
-----------------------
Equation:
int + int = int
0.06*x + 0.095*(9885-x) = 696.35
60x + 95*9885 - 95x = 696350
-------------------------
-35x = -242725
x = $6935 (amt. invested at 6%
------
9885-6935 = $2950 (amt. invested at 9.5%
====================
Cheers,
Stan H.
--------------------